OpenStudy (anonymous):

evaluate the integral of (x^2+x)dx , [-2,0]

OpenStudy (anonymous):

\[=\lim_{n \rightarrow \infty} \frac{ 2 }{ n } \sum_{1}^{n} [(-2+\frac{ 2i }{ n })^2 + (-2+\frac{ 2i }{ n })]\] The answer is 2/3 but I'm getting other numbers.

OpenStudy (anonymous):

\[=\lim_{n \rightarrow \infty}\frac{ 2 }{ n }\sum_{1}^{n}(2-\frac{ 6i }{ n }+\frac{ 4i^2 }{ n^2 })\] \[=\lim_{n \rightarrow \infty}[2n-\frac{ 6 }{ n }(\frac{ n(n+1) }{ 2 })+\frac{ 4 }{ n^2 }(\frac{ n(n+1)(2n+1) }{ 6 })]\] \[=\lim_{n \rightarrow \infty}\frac{ 2 }{ n }(-n-3+\frac{ 4n^2+4n+1 }{ 3n })\] \[=\lim_{n \rightarrow \infty}(-2-3+\frac{ 4n }{ 3 }+\frac{ 4 }{ 3 }+\frac{ 1 }{ 3n })\] \[=\frac{ -11 }{ 3 }\] Help?

OpenStudy (whpalmer4):

I'm puzzled by your method of doing integrals... \[\int\limits_{-2}^{0}(x^2+x) dx = \int\limits_{-2}^{0}x^2 dx + \int\limits_{-2}^{0}x dx\]By the power rule, \[ \int x^2 dx + \int x dx = (\frac{1}{3}x^3+\frac{1}{2}x^2 + C)\] Now evaluate that at \(x = 0\) and \(x = -2\) and subtract the latter from the former.

OpenStudy (anonymous):

Sorry, should've clarified...the problem asks to evaluate the integral using the form of the definition of the integral \[\int\limits_{-2}^{0}f(x)dx=\lim_{n \rightarrow \infty}\sum_{1}^{n}f(x _{i})\Delta x\] where \[\Delta x=\frac{ b-a }{ n } and x _{i}=a+i \Delta x\]

OpenStudy (tkhunny):

You should also be aware of the fact that the curve crosses the x-axis in [-2,0]. It may or may not be important, but you WILL want to remember that, at times.

OpenStudy (anonymous):

I checked on stewart's homework hints and confirmed I'm right up until I get to \[=\lim_{n \rightarrow \infty}\frac{ 2 }{ n } \sum_{1}^{n}[(-2+\frac{ 2i }{ n })^2+(-2+\frac{ 2i }{ n })]\] from here on out it should be algebra and summation distribution but it doesn't give me 2/3

OpenStudy (anonymous):

Hmm..

OpenStudy (anonymous):

check the third step, you should have had -3/n when you distributed n

OpenStudy (anonymous):

oh wait, no that's not it

OpenStudy (anonymous):

when you factored out 2/n, you had -n, when it should be +n

OpenStudy (anonymous):

Woops, 2/n wasn't factored out. It goes after the limit in the step before it - just forgot to write it in.